Make up an equation involving a fractional coefficient and show 2 different methods of solving it.

please and thankyou

Answer by stanbon(75887)

Make up an equation involving a fractional coefficient and show 2 different methods of solving it.
------
Example: (1/2)x = 9
Multiply both sides by 2 to get:
x = 18
---------------------
(1/2)x = 9
Divide both sides by 1/2 to get:
x = 9/(1/2)
x = 9*(2/1) = 18
----------------------
Cheers,
Stan H.
===================